OpenStudy (anonymous):

Can someone help me with exponents? (8th grade level)

OpenStudy (anonymous):

This.

OpenStudy (anonymous):

x^6

OpenStudy (anonymous):

x to the power of 6

OpenStudy (anonymous):

@angelloarce how did you get that answer ?

OpenStudy (anonymous):

i did the math man

OpenStudy (msmikaelar):

you add the exponents at the top and the subtract that from the bottom so 7+4+11 an then 11-5= 6

OpenStudy (triciaal):

when you multiply numbers/variables with the same base add the exponents when you divide numbers or variables with the same base subtract the exponents

OpenStudy (anonymous):

I explained it to you in your other post

OpenStudy (anonymous):

@angelloarce I know that but I'm saying how ?? Like what what strategy did you use ?

OpenStudy (anonymous):

@danaxixi i did exactly what @triciaal said

OpenStudy (anonymous):

To multiply the exponents you add them. To divide them you subtract. You first need to add the exponents in the numerator together, then you will subtract the exponent in the denominator from the numerator. For example: A^4 * A^3= A^7 A^7/A^2= A^5

OpenStudy (anonymous):

@mthompson440 thanks for your help ;)